OpenStudy (isbella):

I have two calculus 3 questions help please 1)Find the curvature k(t) of the curve r(t)=(4sint)I+(4sint)j+(4cost)k? 2)find the arclength of the curve r(t)=<8sqrt(2t),e^(8t),e^(-8t)>,0

OpenStudy (isbella):

I think k(t)=|r^1(t)xr^11(t)|/|r^1(t)|^3 is for the curvature but I keep getting it wrong

OpenStudy (anonymous):

You think \[ \kappa (t) = \frac{\|\mathbf r'(t)\times \mathbf r''(t)\|}{\|\mathbf r'(t)\|^3} \]?

OpenStudy (anonymous):

What did you get for \(\mathbf r'(t)\) and \(\mathbf r''(t)\)?

OpenStudy (isbella):

for r^1(t) I got =4cos(t)I+4cos(t)j-4sin(t)k for r^11(t)=-4sin(t)I-4sin(t)j+4cos(t)k

OpenStudy (anonymous):

that looks right...

OpenStudy (perl):

do you need help?

OpenStudy (isbella):

yes i need help on the second one

OpenStudy (anonymous):

I think the vector in the second problem might supposed to be\[r(t)=<8\sqrt{2}t,e ^{8t},e ^{-8t}>\]

OpenStudy (anonymous):

\[Arclength = \sqrt{(\frac{ dr _{x} }{ dt })^{2}+(\frac{ dr _{y} }{ dt })^{2}+(\frac{ dr _{z} }{ dt })^{2}}\]

OpenStudy (anonymous):

r'(t) = (-4 cos t, -4 cos t, 4 sin t) r''(t) = (4 sin t, 4 sin t, 4 cos t) r' x r'' = (-16, 16, 0). ==> k = ||r' x r''|| / ||r'||^3 k = (16 sqrt(2)) / (4^3 sqrt(1 + cos^2(t))) = sqrt(2)/(4sqrt(1 + cos^2(t))).

OpenStudy (anonymous):

Integrate my previous post over your bounds of t to find the whole arclength